Part A: What is the magnification of an astronomical telescope whose objective lens has a focal length of 71cm and whose eyepiece has a focal length of 2.9cm? Follow the sign conventions. Part B: What is the overall length of the telescope when adjusted for a relaxed eye?
Answer #1

The magnification of an astronomical telescope is m=- \frac{f_o}{f_e}=-71/2.9=-24.48

Over all length of a telescope is l=f_o+f_e=71+2.9=73.9 cm